OpenStudy (danielannx3):

What is the sign of the product (–5)(–3)(–8)(–6)? Positive, because the products (–5)(–3) and (–8)(–6) are negative, and the product of two negative numbers is positive Positive, because the products (–5)(–3) and (–8)(–6) are positive, and the product of two positive numbers is positive Negative, because the products (–5)(–3) and (–8)(–6) are negative, and the product of two negative numbers is negative Negative, because the products (–5)(–3) and (–8)(–6) are positive, and the product of two positive numbers is negative

OpenStudy (danielannx3):

@FaiqRaees

OpenStudy (faiqraees):

Okay so lets observe some cases -1 (sign negative) -1 x -1 = 1 (sign positive) -1 x -1 x -1 = -1 (sign negative) -1 x -1 x -1 x -1 = 1 (sign positive) -1 x -1 x -1 x -1 x -1 = -1 (sign negative) From this can you make an observation relating to the number of negative signs on the left side and and the sign of product?

OpenStudy (danielannx3):

its c

OpenStudy (faiqraees):

Wrong. How about you contemplate on the cases I provided and reach a substantial conclusion?

OpenStudy (danielannx3):

but i dont get it

OpenStudy (faiqraees):

Count the number of negative signs (on the left hand) in each case

OpenStudy (danielannx3):

its a i just notice lol

OpenStudy (danielannx3):

i dont get this ughh

OpenStudy (danielannx3):

The signs are positive because there are 2 groups of 2 negative (negative + negative = positive) (positive + positive = positive)

OpenStudy (danielannx3):

so its a

OpenStudy (faiqraees):

Oh sorry it is A

OpenStudy (faiqraees):

Apologies for inconvience